# Creation of a tweezer array for cold atoms utilizing a generative neural network

### v1

@@ Abstract
Optical tweezers have become essential tools for dynamically manipulating objects, ranging from microspheres or biological molecules to neutral atoms. In this study, we demonstrate the creation of tweezer arrays using a generative neural network, which allows for the trapping of neutral atoms with tunable atom arrays. We have successfully loaded cold strontium atoms into various optical tweezer patterns generated by a spatial light modulator (SLM) integrated with generative models. Our approach shortens the process time to control the SLM wtih minimal time delay, eliminating the need for repeated re-optimization of the hologram for the SLM.
@@ Introduction
Figure 1: Overview of the scheme for producing an atom array based on deep learning To trap ${}^{88}$r atoms, a optical tweezer array is utilized. The Spatial Light Modulator (SLM) is controlled by a generative neural network that is based on the U-net model. This generative model is trained using a hologram dataset consisting of tweezer images and corresponding holograms for the SLM. Various tweezer configurations can be created on demand.
@@
Figure 2: Diagram of the neural network based on the U-net model (a) The architecture of generative neural network accepts a pair of images, followed by the U-net model. (b) Down sampling and Up sampling block architecture. The convolution (Conv) layer is used in down sampling block to increase the channel number while the transpose convolution (TConv) layer is added to the Up-sampling block in order to reconstruct the phase information
@@
As the first step to facilitate the training of the generative model, we introduce a large-scale hologram dataset consisting of 2000 pairs of tweezer images and corresponding holograms for the SLM (Holoeye; PLUTO-2.1-VIS-130). Since the tweezer is usually formed in a far-field regime, we use the traditional Fraunhofer zone method to create the dataset for training our model, resulting in the optical reconstruction on the object plane. The diffraction process in the GS algorithm is calculated according to the Fraunhofer diffraction Hecht (1989).
@@
Figure 4: Various atom arrays produced by the generative neural network. (a) Experimental absorption image of $\sim$20 tweezer sites, where each bright spot corresponds to an ensemble of $\sim$30 ${}^{88}$r atoms. Various lattices including square, triangular and honeycomb geometry are formed by the trained generative neural network. The lattice constant is 15 $\mu$m. Absorption images are recorded with a CMOS camera. (b) With the atomic signal at the center of the tweezer trap, we define the SNR as the ratio between $I_{signal}$ and $I_{noise}$. (c) A comparison of algorithm runtime and tweezer array quality reveals that the generative neural network (green) surpasses the GS algorithm (purple) when it comes to the number of iterations needed. The SNR of the tweezer array is achieved without requiring iteration.
@@
A tweezer array generated by the computer-generated holography (CGH) based on the trained generative neural network is readily usable in a neutral atom experiment Wen et al. (2024). Here we produce a large sample of cold ${}^{88}$r atoms in a narrow-line magneto-optical trap (MOT) in a glass cell (see Fig. 3Wen et al. (2024). To obtain such a narrow-line MOT, a ${}^{88}$r atomic beam generated by an oven in a custom-designed vacuum chamber is slowed down and captured in a broad-line 461 nm two-dimensional MOT. Subsequently, a 461 nm laser beam with low power, referred to as the "pushing beam" exerts a force on the atoms in the 2D-MOT, guiding them into a 461 nm 3D MOT in a separate glass cell as shown in Fig. 3. Finally, we create a 689 nm narrow-line MOT in which a cold gas of 2×$\times 10^{6}$ ${}^{88}$r atoms at the temperature of 2-5 $\mu$K are prepared at the center of the glass cell. The sample preparation process typically takes approximately 1 second Wen et al. (2024). To create a tweezer array, programmable CGHs are generated by the use of phase SLMs (PLUTO-2.1-VIS-130) at the lens’ back focal plane of the objective lens as described in Fig. 3. A 515 nm laser beam of 0.5W is sent to the SLM, resulting in about twenty tweezer traps with the trap depth of $k_{B}\times$1.2 mK where $k_{B}$ is the Boltzmann constant.
@@
With a tweezer array being switched on, a few tens of Sr atoms are trapped in each micro trap, which allows us to examine the quality of the tweezer light pattern generated. To record the number of trapped atoms in each trap, a collimated 461nm laser beam is sent through the Sr atom array, and the optical density (OD) is recorded by the CMOS camera (Dhyana 400BSI V3) Wen et al. (2024). To avoid the unwanted reflections, such as stripes and Newtons rings, we employ a double-exposure scheme. The first exposure is performed while the atoms are trapped in the optical tweezers within 10us while the second reference exposure is performed shortly after 0.5 s as a reference image. The OD is calculated according to $OD=ln\frac{I_{atom}-I_{ref}}{I_{atom}-I_{ref}}$ where $I_{atom}$ is the first image with atoms and $I_{ref}$ is the image without atom.
Fig. 4(a) showcases a variety of arrays generated by the generative neural network. These arrays consist of square, triangular, and hexagonal lattices, with each tweezer trap containing tens of Sr atoms. In order to analyze these arrays, we extract the signal-to-noise ratio (SNR) from the absorption images shown in Fig.ure 4(a). This is done by examining the maximum optical density at the center of each tweezer trap ($I_{signal}$) and comparing it to the background fluctuation ($I_{noise}$), as explained in Fig. 4(b).
The main benefit of using the generative neural network for obtaining CGHs is that this approach overcomes the trade-off between computation time and the quality of the final atom arrays. The generative neural network outperforms the GS algorithm and the SNR of the tweezer array is achieved without requiring iteration. For example, the neural network can generate the POH in 0.16s with the SNR 14dB as shown in Fig. 4(c). In contrast, atom arrays produced by the traditional GS algorithm face the problem of speckle noise due to the initial random phase and amplitude. As a result, several iterations are inevitably required to achieve a comparable SNR, which takes about 6 seconds. In comparing the runtime of the algorithm and the quality of the tweezer array, therefore, it is evident that the generative neural network outperforms the GS algorithm in terms of the required number of iterations.
@@ Conclusion
In summary, we have introduced a method for generating tweezer arrays to trap cold atoms utilizing state-of-the-art generative neural networks. Our findings demonstrate the remarkable capability of the U-net based deep learning approach in successfully producing arrays of Sr atoms. Importantly, this method eliminates the need for time-consuming iterative procedures such as the GS algorithm. Furthermore, the Sr atom array discussed in this work is not limited to its current form but can be easily applied to fermionic isotopes. This would open up exciting possibilities for exploring the SU($N$) physics in a tweezer trap He et al. (2019).

### v2

@@ Abstract
Optical tweezers have become essential tools for dynamically manipulating objects, ranging from microspheres or biological molecules to neutral atoms. In this study, we demonstrate the creation of tweezer arrays using a generative neural network, which allows for the trapping of neurtal atoms with tunable atom arrays. We have successfully loaded cold strontium atoms into various optical tweezer patterns generated by a spatial light modulator (SLM) integrated with generative models. Our approach shortens the process time to control the SLM wtih minimal time delay, eliminating the need for repeated re-optimization of the hologram for the SLM.
@@ Introduction
Figure 1: Overview of the scheme for producing an atom array based on deep learning To trap 88Sr atoms, a tweezer array is utilized. The Spatial Light Modulator (SLM) is controlled by a generative neural network that is based on the U-net model. This generative model is trained using the tweezer dataset. The high level of control provided by this setup allows us to create precise atomic configurations.
@@
Figure 2: Diagram of the neural network based on the U-net model (a) The architecture of generative neural network accepts a pair of images, followed by the U-net model. (b) Down sampling and Up sampling block architecture. The convolution (Conv) layer is used in down sampling block to increase the channel number while the transpose convolution (TConv) layer is added to the Upsampling block in order to reconstruct the phase information.
@@
As the first step to facilitate the training of the generative model, we introduce a large-scale hologram dataset consisting of 2000 pairs of tweezer images and corresponding holograms for the SLM (Holoeye; PLUTO-2.1-VIS-130). Since the tweezer is usually formed in a far-field regime, we use the traditional Fraunhofer zone method to create the dataset for training our model, resulting in the optical reconstruction on the object plane. The diffraction process in the GS algothrim is calculated according to the Fraunhofer diffraction Hecht (1989).
@@
Figure 4: Various atom arrays produced by the generative neural network. (a) Experimental absorption image of a $\sim$20 tweezer site, where each bright spot corresponds to an ensemble of $\sim$30 88Sr atoms. Various lattices including square, triangular and honeycomb geometry are formed with the lattice constant of 15 $\mu$m. Absorption images are recorded with a CMOS camera. (b) With the atomic signal at the center of the tweezer trap, we define the SNR as the ratio between $I_{signal}$ and $I_{noise}$. (c) Evaluation of algorithm runtime and tweezer array quality shows that the generative neural network outperforms the GS algorithm. The signal-to-noise ratio (SNR) of the tweezer array is achieved without requiring iteration. The error bar represents the standard deviation.
@@
A tweezer array generated by the computer-generated holography (CGH) based on the trained generative neural network is readily usable in a neutral atom experiment Wen et al. (2024). Here we produce a large sample of cold 88Sr atoms in a narrow-line magneto-optical trap (MOT) in a glass cell (see Fig. 3Wen et al. (2024); Yan et al. (2024). To obtain such a narrow-line MOT, a 88Sr atomic beam generated by an oven in a custom-designed vacuum chamber is slowed down and captured in a broad-line 461 nm two-dimensional MOT. Subsequently, a 461 nm laser beam with low power, referred to as the "pushing beam" exerts a force on the atoms in the 2D-MOT, guiding them into a 461 nm 3D MOT in a separate glass cell as shown in Fig. 3. Finally, we create a 689 nm narrow-line MOT in which a cold gas of 2×$\times 10^{6}$ 88Sr atoms at the temperature of 2-5 $\mu$K are prepared at the center of the glass cell. The sample preparation process typically takes approximately 1 second Wen et al. (2024). To create a tweezer array, programmable CGHs are generated by the use of phase SLMs (PLUTO-2.1-VIS-130) at the lens’ back focal plane of the objective lens as described in Fig. 3. A 515 nm laser beam of 0.5W is sent to the SLM, resulting in about twenty tweezer traps with the trap depth of $k_{B}\times$1.2 mK where $k_{B}$ is the Boltzmann constant.
@@
With a tweezer array being switched on, a few tens of Sr atoms are trapped in each micro trap, which allows us to examine the quality of the tweezer light pattern generated. To record the number of trapped atoms in each trap, a collimated 461nm laser beam is sent through the Sr atom array, and the optical density (OD) is recorded by the CMOS camera (Dhyana 400BSI V3) Wen et al. (2024). To avoid the unwanted reflections, such as stripes and Newtons rings, we employ a double-exposure scheme. The first exposure is performed while the atoms are trapped in the optical tweezers within 10us while the second reference exposure is performed shortly after 0.5 s as a reference image. The OD is calculated according to $OD=ln\frac{I_{atom}-I_{ref}}{I_{atom}+I_{ref}}$ where $I_{atom}$ is the first image with atoms and $I_{ref}$ is the image without atom.
Fig. 4(a) showcases a variety of arrays generated by the generative neural network. These arrays consist of square, triangular, and hexagonal lattices, with each tweezer trap containing tens of Sr atoms. In order to analyze these arrays, we extract the signal-to-noise ratio (SNR) from the absorption images shown in Fig. 4(a). This is done by examining the maximum optical density at the center of each tweezer trap ($I_{signal}$) and comparing it to the background fluctuation ($I_{noise}$), as explained in Fig. 4(b).
The main benefit of using the generative neural network for obtaining CGHs is that this approach overcomes the trade-off between computation time and the quality of the final atom arrays. The generative neural network outperforms the GS algorithm and the SNR of the tweezer array is achieved without requiring iteration. For example, the neural network can generate the POH in 0.16s with the SNR 14dB as shown in Fig. 4(c). In contrast, atom arrays produced by the traditional GS algorithm face the problem of speckle noise due to the initial random phase and amplitude. As a result, several iterations are inevitably required to achieve a comparable SNR, which takes about 6 seconds.
@@ Conclusion
In summary, we have introduced a method for generating tweezer arrays to trap cold atoms using state-of-the-art generative neural networks. Our findings demonstrate the remarkable capability of the U-net based deep learning approach in successfully producing arrays of Sr atoms. Importantly, this method eliminates the need for time-consuming iterative procedures such as the GS algorithm. Furthermore, the Sr atom array discussed in this work is not limited to its current form but can be easily applied to fermionic isotopes. This would open up exciting possibilities for exploring the SU($N$) physics in a tweezer trap He et al. (2019).
Data Availability
The data that supports the findings of this study are available within the article.